A New York City daily newspaper called “Manhattan Today” charges an annual subscription fee of $135. Customers prepay their subscriptions and receive 260 issues over the year. To attract more subscribers, the company offered new subscribers the ability to pay $130 for an annual subscription that also would include a coupon to receive a 40% discount on a one-hour ride through Central Park in a horse-drawn carriage. The list price of a carriage ride is $125 per hour. The company estimates that approximately 30% of the coupons will be redeemed.

Required:
1. How much revenue should Manhattan Today recognize upon receipt of the $130 subscription price?
2. How many performance obligations exist in this contract?
3. Prepare the journal entry to recognize sale of 10 new subscriptions, clearly identifying the revenue or deferred revenue associated with each performance obligation.
Answer #1
1
Revenue = $0
2
Number of performance obligations = 2
3
General Journal Debit Credit
Cash 1300
      Deferred revenue—subscription 1170 =1300*90%
      Deferred revenue—discount coupon 130 =1300*10%
Workings:
Stand­alone selling price of coupon 15 =125*40%*30%
Stand­alone selling price of subscription 135
Total of stand­alone prices 150
Percentage allocated to—subscription 90% =135/150
Percentage allocated to—discount coupon 10% =15/150